### v1

# Dance Any Beat: Blending Beats with Visuals in Dance Video Generation

Abstract
The task of generating dance from music is crucial, yet current methods, which mainly produce joint sequences, lead to outputs that lack intuitiveness and complicate data collection due to the necessity for precise joint annotations. We introduce a Dance Any Beat Diffusion model, namely DabFusion, that employs music as a conditional input to directly create dance videos from still images, utilizing conditional image-to-video generation principles. This approach pioneers the use of music as a conditioning factor in image-to-video synthesis. Our method unfolds in two stages: training an auto-encoder to predict latent optical flow between reference and driving frames, eliminating the need for joint annotation, and training a U-Net-based diffusion model to produce these latent optical flows guided by music rhythm encoded by CLAP. Although capable of producing high-quality dance videos, the baseline model struggles with rhythm alignment. We enhance the model by adding beat information, improving synchronization. We introduce a 2D motion-music alignment score (2D-MM Align) for quantitative assessment. Evaluated on the AIST++ dataset, our enhanced model shows marked improvements in 2D-MM Align score and established metrics. Video results can be found on our project page: https://DabFusion.github.io.
Figure 1: We present a diffusion-based framework DabFusion to dance any beat given any image. With various music styles of AIST++ dataset, our model predicts matching future video frames (128 $\times$ 128) starting from the same reference image.
1 Introduction
Music-to-dance generation offers significant potential for automating choreography, benefiting various real-world applications. Existing methods [26, 27, 47, 35] focus on generating joint sequences, which, despite being effective, often produce outputs that are less intuitive for users and require precise joint annotations during data collection. This complicates the process. The field of conditional image-to-video (I2V) generation—where videos are produced from a single still image guided by textual descriptions [54, 39, 11], driving videos [6, 65, 55], or pose sequences [38, 45, 44, 68]—offers a way to address these challenges. However, using music as a driving condition remains underexplored, presenting an opportunity for innovation. In this paper, we propose the Dance Any Beat Diffusion model (DabFusion), which generates dance videos from still images and music, enhancing intuitiveness and eliminating the need for joint annotations. DabFusion also pioneers the use of music as a condition in image-to-video synthesis.
DabFusion is highly versatile, capable of generating dance videos using the AIST++ dataset and animating individuals in new scenarios. As shown in Figure 1, we animate previously unseen individuals by merging their images with a background from the AIST++ dataset, bringing them to life through dance. This capability makes DabFusion practical for real-world applications, such as aiding dancers and choreographers in learning and practice, fostering engagement on social media, and enabling users without formal dance training to create personalized dance content. When generating dance videos using the AIST++ dataset, as illustrated in Figure 2, DabFusion excels in producing varied dance styles with different dancers from multiple perspectives and diverse initial poses. This flexibility enhances the intuitiveness and applicability of the generated content.
Inspired by recent I2V works [32, 44, 57], our model’s training process is structured into two phases. First, a latent flow auto-encoder is trained unsupervisedly to estimate the latent optical flow between reference and driving frames in a video, aiding in warping the reference frame to generate movement. Then, a U-Net based diffusion model generates latent flows guided by music and a starting image. For music information extraction, we use CLAP [59], the largest audio representation foundation model, to facilitate high-quality, music-aligned dance video generation.
A critical aspect of dance video evaluation is synchronization between the dancer’s movements and the music’s beat, an area needing improvement in our initial model. To address this, we integrate Librosa [30], a tool for audio signal analysis and beat extraction, with CLAP’s output. This enhancement improves alignment between dance poses and music beats, offering a more coherent and engaging viewing experience. To quantitatively assess this alignment, we introduce the 2D motion-music alignment (2D-MM Align) score, inspired by the Beat Alignment Score [27] for 3D motion-music correlation. This metric evaluates synchronization between motion and music in 2D scenarios, demonstrating the effectiveness of our approach in producing rhythmically aligned dance videos.
Figure 2: Exemplar videos generated from our DabFusion. Taking first image as starting frame and the unique music clip as guiding dance style, our framework is capable of generating varied styles of dance videos featuring different dancers from multiple perspectives with diverse initial poses and positions.
Our contribution can be summarized as follows:
  • We introduce a novel approach capable of directly generating dance videos from music, significantly enhancing the intuitiveness of the generated content and its applicability in real-world settings.
  • Our research broadens the horizon of conditional image-to-video generation by exploring music as a novel conditional input. We investigate the extraction of meaningful information from music through the integration of large-scale foundation models and established signal-processing methodologies.
  • We establish a baseline and propose a beat alignment module for this innovative task, supported by extensive evaluations and the introduction of a new evaluation metric to assess motion-music alignment.
2 Related Works
Generating dance sequences from music uniquely intersects motion synthesis [1, 4, 12, 14] and music interpretation [58, 30, 59], aiming to create choreographed movements synchronized with input music. This extends beyond traditional motion synthesis, as choreographed movements are complex to animate. Early research [43, 69, 9] focused on producing 2D dance sequences due to the availability of online dance videos and advances in 2D human pose estimation [5]. However, 2D predictions lack expressiveness and applicability, prompting a shift to 3D dance generation. Recent methods leverage LSTMs [49, 2, 20, 61], GANs [25, 48, 10, 21], and Transformers [17, 27, 47, 35, 64] for 3D motion generation. The AIST++ dataset [27], a rich compilation of 3D motion data, has significantly advanced the field. Notable models include FACT [27], a full-attention-based cross-modal transformer using sequence-to-sequence learning for lifelike 3D dance sequences, and Bailando [47], which combines a pose VQ-VAE with a Motion GPT for temporal coherence via actor-critic learning. While existing research focuses on generating motion sequences from music, Ren et al. [37] synthesize videos using these sequences. Our work bypasses intermediary joint sequence generation, directly creating dance videos from music.
Image-to-Video (I2V) generation [3, 8, 16, 29, 34, 54, 60, 62, 66, 55] transforms static images into dynamic video sequences. This domain relates closely to single-image video prediction [22, 28], which infuses still images with temporal continuity to create believable motion. A specialized subfield, conditional image-to-video generation [3, 8, 16, 29, 54, 62, 55], incorporates guiding conditions or cues into the video generation process, enriching the content with specific characteristics or actions. For example, Wang et al. [54] developed a conditional GAN architecture to create human videos from static images, conditioned on labels describing facial expressions or actions. Yang et al. [62] introduced a pose-guided approach that extracts poses from an image and uses GANs to generate a sequence of poses and then video frames, translating static postures into animated sequences. Building on these, Wang et al. [55] unveiled the Latent Image Animator (LIA), which animates still images into videos under the guidance of an auxiliary video, leveraging a self-supervised auto-encoder to navigate latent space for animation. However, using audio, particularly music, as a conditioning factor in I2V remains largely unexplored. Introducing musical elements as conditions for video generation could open new research avenues and applications, suggesting an exciting direction for future work in I2V technology.
3 Method
Our methodology aims to generate latent optical flows guided by musical inputs. Recent works [31, 44, 52, 53, 32] in motion transfer have shown the effectiveness of using latent optical flow for warping images. This approach is more resource-efficient, requiring less computational power and time compared to high-dimensional pixel or latent feature spaces. Diffusion models are chosen for their exceptional quality and robust controllability. Our methodology begins by training an auto-encoder to discern optical flow between video frames. This trained auto-encoder then aids in training the diffusion model to generate latent flows. We extract musical information using CLAP to encode the music. Our enhanced model also incorporates beat information for better representation.
Figure 3: Training of latent flow auto-encoder. The flow predictor learns to estimate the latent flow $f$ and occlusion map $m$ between the reference frame $x_{ref}$ and driving frame $x_{dri}$. The image encoder encodes $x_{ref}$ into a latent representation $z$, $f$ and $m$ are utilized to manipulate $z$ into $\tilde{z}$ which is then decoded by an image decoder to generate a output image $\hat{x}_{out}$. The objective of the training is to minimize the disparity between $x_{dri}$ and $\hat{x}_{out}$.
3.1 Latent Flow Estimation
The objective of this phase is the training of an auto-encoder capable of accurately capturing and modeling motion between video frames within a latent space. As illustrated in Figure 3, the architectural components of the model comprise an image encoder, a flow predictor, and an image decoder. During the training process, two frames are randomly selected from a single video to serve as a reference frame $x_{ref}$ and a driving frame $x_{dri}$, both sharing dimensions of $H_{x}\times W_{x}\times 3$. The image encoder processes $x_{ref}$, converting it into a compact latent representation, $z$, with dimensions $H_{z}\times W_{z}\times C_{z}$. Subsequently, similar to [44, 53], the flow predictor, receiving both $x_{ref}$ and $x_{dri}$ as inputs, computes the backward latent optical flow, $f$, and an occlusion map, $m$, to denote the transformation between these frames. The estimated latent flow, $f$, maintains the spatial dimensions of $z$ and incorporates two channels to articulate the horizontal and vertical displacement across frames. Using backward flow estimation, the warping of $z$ using $f$ is realized through an efficient differentiable bilinear sampling [18]. Concurrently, the occlusion map, sized $H_{z}\times W_{z}\times 1$, facilitates the reconstruction of areas within $z$ that become obscured or revealed due to movement. Values within this map range from 0 to 1, where 1 signifies unoccluded regions, and 0 denotes complete occlusion. The warped latent map, $\tilde{z}$, is derived through the following equation:
$\tilde{z}\ =\ m\ \odot f_{w}(z\ ,\ f),$ (1)
where $f_{w}(\cdot\ ,\cdot)$ denotes the back-warping operation and $\odot$ is the element-wise multiplication. Finally, the image decoder takes $\tilde{z}$ as input, reconstructing visible portions while simultaneously inpainting occluded regions to produce the output image $\hat{x}_{out}$. The training objective is to minimize a reconstruction loss that quantifies the difference between $\hat{x}_{out}$ and $x_{dri}$. This loss employs the perceptual loss proposed by Johnson et al. [19], which utilizes features extracted by a pre-trained VGG-19 network [46]. Formally, the reconstruction loss is expressed as:
$\mathcal{L}_{rec}\ =\ \sum_{i=1}^{N}\left|V_{i}(\hat{x}_{out})-V_{i}(x_{dri})\right|,$ (2)
where $V_{i}(\cdot)$ extracts the ith channel features from a specified layer of VGG-19, and N is the total number of feature channels within that layer.
Figure 4: Inference process of DabFusion. DabFusion incorporates noise input along with image embedding $z_{0}$ and music embedding $e$ as conditions. Following the denoising stage of the diffusion model, we obtain $h_{1}^{N}$, comprising a concatenated sequence of latent flow and corresponding occlusion maps. $h_{1}^{N}$ is utilized to transform $z_{0}$ into a new sequence of latent maps, denoted as $\tilde{z}_{1}^{N}$, which is subsequently decoded to produce an image sequence.
3.2 Latent Flow Generation
We use diffusion models to generate latent flow. These generative models reverse a diffusion process, which gradually turns data into Gaussian noise. The reverse process then restores this noise back to the original data distribution. Mathematically, the diffusion process is a Markov chain that adds noise to the data over $T$ steps.The process starts with the original data $x_{0}$ and ends with a sample $x_{T}$ that resembles Gaussian noise. The transition from $x_{t-1}$ to $x_{t}$ can be defined by the following equation:
$x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t},$ (3)
where $\alpha_{t}$ is a variance schedule that determines the amount of noise added at each step, and $\epsilon_{t}$ is sampled from a standard Gaussian distribution $\mathcal{N}(0,\mathcal{I})$. The reverse diffusion process seeks to reconstruct the original data from noise by learning the conditional distribution $p(x_{t-1}|x_{t})$. This is achieved by training a denoising model to estimate the parameters of the Gaussian distribution of $x_{t-1}$ given $x_{t}$. The model outputs parameters $\mu_{\theta}(x_{t},t)$ and $\sigma_{\theta}(x_{t},t)$, which define the Gaussian distribution from which $x_{t-1}$ is sampled:
$x_{t-1}=\mu_{\theta}(x_{t},t)+\sigma_{\theta}(x_{t},t)\cdot\epsilon,$ (4)
where $\epsilon$ is again sampled from a standard Gaussian distribution. The model is trained to minimize the difference between the generated and actual data distributions, with an objective function often approximating the negative log-likelihood of the data under the model. A practical approach for training diffusion models is employing the Mean Squared Error (MSE) between the predicted and actual noise used in data corruption at each diffusion step. Given the diffusion process, the goal during training is to predict the noise $\epsilon$ that was added to the original data at each step. The loss function for a single training step can be formulated as:
$\mathcal{L}=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right],$ (5)
where $\epsilon_{\theta}(x_{t},t)$ is the predicted noise by the model given the noised data $x_{t}$ at time step $t$. $\epsilon$ is the actual noise added to the data $x_{0}$ to obtain $x_{t}$.
Our approach utilizes a 3D U-Net [7] as the denoising mode. And the training of the model is in conjunction with the trained image encoder and flow predictor. Given an input video $x_{0}^{N}=\{x_{0},x_{1},...,x_{N}\}$ and its corresponding music $m$. We first use the flow predictor to estimate latent flow sequence $f_{1}^{N}=\{f_{1},...,f_{N}\}$ and occlusion map sequence $m_{1}^{N}=\{m_{1},...,m_{N}\}$ between starting frame $x_{0}$ and remaining frames $\{x_{1},...,x_{N}\}$. The size of $f_{1}^{N}$ is $N\times H_{z}\times W_{z}\times 2$ and the size of $m_{1}^{N}$ is $N\times H_{z}\times W_{z}\times 1$. Subsequently, we concatenate these two sequences along the last dimension to get $a_{0}=[f_{1}^{N},m_{1}^{N}]$ which has the size $N\times H_{z}\times W_{z}\times 3$. Then, we perform diffusion process which gradually add 3D Gaussian noise to $a_{0}$ to map it to a a standard Gaussian noise. This is then reversed, using the image encoder to map the initial frame $x_{0}$ to latent space $z_{0}$ and incorporating music embeddings e for conditioning the denoising model. We will discuss how to encode music in Sec. 3.3. The updated loss function, accounting for the conditioning on $z_{0}$ and e, is given by:
$\mathcal{L}=\mathbb{E}_{t,a_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(a_{t},t,z_{0},e)\|^{2}\right].$ (6)
The process of inference is shown in Figure 4 , given an initial image $x_{0}$ and a piece of music $m$, the image is first encoded into a latent space representation $z_{0}$, while the music is transformed into an embedding $e$. Subsequently, a volume of randomly sampled Gaussian noise, with dimensions $N\times H_{z}\times W_{z}\times 3$, undergoes a gradual denoising process by the U-Net. This process yields $h_{1}^{N}$ which is the concatenation of a latent flow sequence $f_{1}^{N}$ and corresponding occlusion map sequence $m_{1}^{N}$. $z_{0}$ is then warped by each latent flow in the latent flow sequence and corresponding occlusion map in the occlusion map sequence to generate a new latent map sequence $\tilde{z}_{1}^{N}$. These maps are sequentially fed into the image decoder, which synthesizes the frames of the new video.
3.3 Music Encoding
Extracting meaningful information from music is crucial to our task. We use CLAP, developed by Wu et al. [59], for music encoding. CLAP is a state-of-the-art audio representation model trained on a large dataset of text-audio pairs, learning a joint text-audio embedding space through contrastive learning. Our choice of CLAP is strategic, aiming to decode dance style information from musical cues using text as a bridge for music-to-dance style translation. While CLAP excels at encoding musical style, it struggles with detecting rhythmic patterns. To address this, we integrate beat information extracted via Librosa with CLAP’s output, creating a composite music embedding. This enhanced approach improves synchronization between motion and beat compared to our baseline model.
4 Experiments
4.1 Dataset and Implementation Details
Our model is trained and evaluated using the AIST++ [27] dataset, a leading resource in 3D human dance. AIST++ is recognized as the most comprehensive dataset in its domain, encompassing 10 dance genres. Each genre features 6 unique pieces of music, with each piece accompanying multiple videos, totaling 12,670 videos. These videos showcase a range of choreographies from basic to advanced, offering a wide spectrum of movements for detailed analysis. The pivotal advantage of AIST++ for our research is its meticulous synchronization of music and dance movements, which is crucial for music-driven dance generation and analysis.
We organize our dataset by music pieces, allocating 10,564 videos and 50 music tracks to the training set, and 2,106 videos with 10 tracks to the test set. Each video is segmented into frames, resized to $128\times 128$ pixels. In the latent flow estimation phase, we adopt the architecture from [19] for both our image encoder and decoder. The flow predictor follows [44]. The model is trained over 150 epochs with a batch size of 100, using the Adam optimizer [23], starting with a learning rate of $2\times 10^{-4}$, decreased by a factor of 0.1 after epochs 60, 90, and 120. For the denoising model’s training, we employ a conditional 3D U-Net architecture from [13], featuring four down-sampling and up-sampling 3D convolutional blocks. Time step $t$ is encoded by the sinusoidal position embedding [51]. The training spanned 250 epochs, starting with the same initial learning rate as the first stage, and adjusted at epochs 100, 150, and 200. We set $T$ = 1000 and apply a cosine noise schedule [33] and dynamic thresholding [41] at the 90% during sampling.
Figure 5: Video quality comparison between our models. We use the same starting image and music piece to generate videos with both our models, and select three frames from the same position.
4.2 Evaluation Metrics
To assess the quality of generated videos, we use the Fréchet Video Distance (FVD) [50]. For quantitative assessment of image-level quality, we employ SSIM [56], PSNR [15], and LPIPS [67]. Additionally, we use the CLIP-Score (CS) [42] to measure the cosine similarity between CLIP embeddings of the dance video and Wav2CLIP [58] embeddings of the music, evaluating the coherence between the dance video and the music. To evaluate synchronization between the dancer’s movements and the music beat, we introduce the 2D motion-music alignment score (2D-MM Align). This metric, inspired by the Beat Alignment Score [27] for 3D scenes, adapts the concept for 2D environments. The alignment score is calculated as the average distance between each kinematic beat and its closest music beat. 2D-MM Align is defined as:
[…]
where $F^{y}=\{f_{j}^{y}\}$ is the music beats, $F^{x}=\{f_{i}^{x}\}$ is the kinematic beats and $\sigma$ is a normalized parameter. Additionally, we incorporate the Audio-video Alignment Score (AV-Align) [63] to quantify audio-video synchronization. This metric analyzes optical flow in the video stream and audio peaks in the audio stream and assesses their temporal alignment within a three-frame window.
Figure 6: Exemplar videos generated using our DabFusion. Showcasing dance videos with unseen individuals in various poses. Following preprocessing of the input image, DabFusion demonstrates its ability to produce dance videos of high quality.
[…]
MM-Diffusion 180.31 0.025 34.81 0.951
[…]
Table 2: Quantitative results across different angles. C01 is in front of the dancer, C02 is in the upper right of the dancer, C03 is in the right of the dancer, C09 is in front of the dancer but closer than C01.
[…]
MM-Diffusion C01 176.93 0.019 35.52 0.949
C01 158.16 0.029 26.41 0.963
[…]
C03 158.56 0.036 26.30 0.959
C09 195.83 0.037 25.09 0.953
C01 158.02 0.027 26.33 0.964
[…]
C03 157.94 0.036 26.29 0.960
C09 195.17 0.035 25.28 0.956
4.3 Result Analysis
[…]
A notable characteristic of the AIST++ dance dataset is its use of nine cameras C01-C09 to capture dance performances, offering multiple perspectives. We analyze four specific angles, summarized in Table 2, to assess their impact on video quality. MM-Diffusion, trained on a subset of the AIST++ dataset using only data from the C01 camera, is used to generate an equivalent number of videos for direct comparison. Our models outperform MM-Diffusion in terms of FVD and SSIM scores, especially for camera position C01, demonstrating superior video quality. Analysis of scores among different cameras indicates that direct camera angles, such as C01 and C03, where the cameras point straight at the dancer, result in higher-quality video generation compared to oblique angles like C02. Historically, I2V research has not emphasized the impact of camera positioning. Our findings reveal that camera placement significantly influences the quality of resulting videos, underscoring its importance in future I2V research. Additionally, we explored the impact of camera distance by comparing videos from C01 and C09, which share the same angle but differ in their proximity to the dancer. The findings suggest that cameras positioned further from the dancer tend to yield better results.
Table 3: Alignment results between our models: video-audio synchrony and motion-beat alignment.
[…]
GT 0.153 0.223 0.152
[…]
Table 4: Alignment results based on dance styles.
[…]
GT 0.138 0.231 0.226
[…]
GT 0.063 0.198 0.101
[…]
GT 0.121 0.218 0.134
[…]
Alignment Evaluation. In Table 3, we assess the synchronicity between music-video and music-motion pairings. The results indicate that our enhanced model surpasses the baseline in all evaluation metrics, demonstrating the positive impact of beat information on both video-audio and motion-beat alignment. Figure 5 illustrates the differences, showing that frames generated by our enhanced model more closely resemble the ground truth videos. We also evaluate the results according to dance style categories corresponding to the music. These assessments, detailed in Table 4, exhibit significant disparities in alignment scores among various dance genres. Notably, Ballet Jazz yields the lowest alignment scores, attributed to its challenging choreography compared to other dance styles. Consequently, the alignment scores reflect the varying levels of difficulty inherent to each dance genre.
4.4 Choreograph Anyone
DabFusion can choreograph individuals whether they belong to the AIST++ dataset or not, requiring several preprocessing steps. First, we use YOLO [36] for object detection and the Segment Anything Model [24] for segmentation to identify and segment the person from unseen scenarios. We then substitute the individual in the image from the AIST++ dataset with the newly segmented person, creating a combined image. This fused image is set as the starting frame to generate dance videos with musical conditions. We randomly select images featuring diverse individuals in various poses to create dance videos. As illustrated in Figure 6, preprocessing of the input image results in a new image compatible with DabFusion for dance video generation.
Table 5: Ablation study for result comparison between 40-frame video and 80-frame video. The results are from our enhanced model.
[…]
40 frames 0.142 0.215 0.148 193.98
80 frames 0.153 0.219 0.179 194.84
Table 6: Ablation study for result comparison between Music embedding and text embedding. The results are from our baseline model.
[…]
text 0.136 0.181 0.126 200.54
music 0.138 0.202 0.141 194.44
Table 7: Ablation study for result comparison about arbitrary length video generation. The results are from our enhanced model. We use the last frame of the first generated videos as the starting frame of next video to generate longer videos.
[…]
40 frames 0.142 0.215 0.148 193.98
80 frames 0.145 0.203 0.139 210.86
120 frames 0.131 0.197 0.121 256.32
160 frames 0.123 0.179 0.101 323.72
4.5 Ablation Study
Influence of Video Length. We observe that video length impacts benchmark performance. Our models are trained to generate videos of 40 and 80 frames. Due to training time and computational resource considerations, we designate the 40-frame videos as the baseline. This comparison is detailed in Table 5. The results indicate comparable video quality between the two lengths; however, the 80-frame videos demonstrate superior performance in alignment evaluations. This improvement is primarily attributed to the extended music duration, providing more information for enhancing alignment accuracy.
Text Embedding vs. Music Embedding. Given that CLAP is trained on text-audio pairs, we aim to uncover the type of information it extracts from music, hypothesizing that CLAP primarily identifies dance style cues. To test this, we conduct an experiment using our baseline model, focusing solely on the embedding from CLAP. We substitute the music embedding with one derived from CLAP using the text description of the dance style corresponding to the music. The findings, presented in Table 6, indicate that the impact of different embeddings on FVD and CS scores is minimal. However, videos generated with text-based embeddings exhibit inferior motion-beat alignment. This suggests that CLAP extracts more than just dance style information; it also obtains details crucial for enhancing motion-beat synchronization.
Arbitrary-Length Video Generation. Given the specialty of our task, our model theoretically can generate videos corresponding to any length of music. By using the final frame of one video sequence as the initial frame of the subsequent sequence, we can create videos of arbitrary length. We conducted experiments with four varying lengths of videos generated sequentially using our enhanced model; the results are presented in Table 7. As video length increases, quality tends to drop rapidly. This decline is primarily because the final frames generated by the model lack the fidelity of actual video frames, becoming more apparent with longer videos. Additionally, alignment scores diminish due to this quality degradation. For instance, employing a pose estimation algorithm to determine keypoints for calculating the 2D-MM Alignment score becomes less accurate with lower video quality, adversely affecting the alignment score.
5 Conclusion
In this paper, we explore the intersection of two distinct research areas: conditional image-to-video generation and music-to-dance generation. We introduce a novel approach, DabFusion, that directly generates dance videos from a single image and accompanying music. This research investigates the role of music as a conditional input in image-to-video synthesis and sets a foundational benchmark for direct dance video generation in response to musical cues. We further examine the extraction of crucial information for enhancing motion-music beat alignment, leveraging both the large-scale foundational model CLAP and the signal-processing tool Librosa. Through extensive experimentation, we have established a robust baseline for this novel task.