### v2

# Dance Any Beat: Blending Beats with Visuals in Dance Video Generation

Abstract
Automated choreography advances by generating dance from music. Current methods create skeleton keypoint sequences, not full dance videos, and cannot make specific individuals dance, limiting their real-world use. These methods also need precise keypoint annotations, making data collection difficult and restricting the use of self-made video datasets. To overcome these challenges, we introduce a novel task: generating dance videos directly from images of individuals guided by music. This task enables the dance generation of specific individuals without requiring keypoint annotations, making it more versatile and applicable to various situations. Our solution, the Dance Any Beat Diffusion model (DabFusion), utilizes a reference image and a music piece to generate dance videos featuring various dance types and choreographies. The music is analyzed by our specially designed music encoder, which identifies essential features including dance style, movement, and rhythm. DabFusion excels in generating dance videos not only for individuals in the training dataset but also for any previously unseen person. This versatility stems from its approach of generating latent optical flow, which contains all necessary motion information to animate any person in the image. We evaluate DabFusion’s performance using the AIST++ dataset, focusing on video quality, audio-video synchronization, and motion-music alignment. We propose a 2D Motion-Music Alignment Score (2D-MM Align), which builds on the Beat Alignment Score to more effectively evaluate motion-music alignment for this new task. Experiments show that our DabFusion establishes a solid baseline for this innovative task. Video results can be found on our project page: https://DabFusion.github.io.
Figure 1: We present a diffusion-based framework DabFusion to dance any beat given any image. With various music styles of AIST++ dataset, our model predicts matching future video frames (128 $\times$ 128) starting from the same reference image.
1 Introduction
Generating dance from music has great potential for automating choreography and enhancing various applications. Current methods [25, 26, 46, 34] primarily produce keypoint sequences which, while effective, often yield outputs that are not as user-friendly and cannot animate specific individuals, limiting their use in personalized scenarios. These methods also rely heavily on precisely annotated keypoints, complicating data collection and limiting their application to non-standard or self-made video datasets. To address these limitations, we introduce a new and innovative task: generating dance videos directly from images of individuals, guided by accompanying music. This allows for personalized dance videos featuring specific people without needing keypoint annotations, making it more versatile for different scenarios. Our solution, named the Dance Any Beat Diffusion model (DabFusion), takes a reference image paired with a selected piece of music as input to create dance videos featuring various dance types and choreographies. For extracting information from music, we develop a novel music encoder that captures essential features including dance style, movement, and rhythm, which are critical for this task. To accurately extract dance styles, we fine-tune the CLAP [58] model, a text-audio pretraining foundation model, on the AIST++ [26] dataset, which serves as the training dataset for our task. For movement information, we fine-tune the Wav2CLIP [57] model, a visual-audio pretraining foundation model, on the AIST++ dataset. The synchronization of dance movements with the music’s rhythm is essential for realistic dance generation. To achieve this, we employ Librosa [29], a tool for audio signal analysis, to extract beat information from the music. By integrating these elements, our music encoder can generate a comprehensive representation of the music.
DabFusion effectively generates dance videos not just for individuals included in the AIST++ dataset, but also for people it has never encountered before. As shown in Fig. 1, we animate previously unseen individuals by merging their images with a background from the AIST++ dataset, bringing them to life through dance. DabFusion can choreograph dance sequences for previously unseen individuals to various music pieces, producing distinctly different videos. This diversity in output highlights the robust information extraction capability of our music encoder. When generating dance videos using the AIST++ dataset, as illustrated in Fig. 2, DabFusion excels in producing varied dance styles with different dancers from multiple perspectives and diverse initial poses. The capability of DabFusion to generate various dance forms enables it to serve a broad audience, ranging from professional choreographers seeking inspiration to casual users wanting to see themselves or others dance to their favorite tunes. Additionally, this technology can be utilized in entertainment applications, such as music videos and virtual concerts, where personalized and accurately synchronized dance performances can greatly enhance the viewer experience. The versatility of DabFusion arises from its method of generating latent optical flow, which captures all essential motion information needed to animate any person in the image. This technique is widely adopted in the field of image-to-video (I2V) generation [53, 38, 10, 5, 64, 54, 37, 44, 43, 67]. Inspired by recent I2V works [31, 43, 56], our model’s training process is structured into two phases. First, a latent flow auto-encoder is trained unsupervisedly to estimate the latent optical flow between reference and driving frames in a video, aiding in warping the reference frame to generate movement. Then, a U-Net based diffusion model generates latent flows guided by music and a starting image. We conduct a thorough evaluation of DabFusion’s performance using the AIST++ dataset, with particular focus on video quality, audio-video synchronization, and the alignment between motion and music. To more effectively assess motion-music alignment—a crucial aspect of this new task—we introduce a 2D motion-music alignment (2D-MM Align) score, inspired by the Beat Alignment Score [26] for 3D motion-music correlation. This metric evaluates synchronization between motion and music in 2D scenarios, demonstrating the effectiveness of our approach in producing rhythmically aligned dance videos. Our experimental results demonstrate that DabFusion establishes a robust baseline for this innovative task, indicating its potential as a pioneering tool in the field of dance video synthesis.
Our contributions can be summarized as follows:
  • We introduce a novel task that directly generates dance videos from music, significantly enhancing the intuitiveness of the generated content and its applicability in real-world settings.
  • We propose DabFusion, a solution for the new task. DabFusion is capable of creating dance sequences not only for individuals present in the training dataset but also for person it has never previously encountered.
  • We propose a novel music encoder designed to generate a comprehensive representation of music, capturing key elements including dance style, movement, and rhythm.
  • We establish a solid baseline for the new task, supported by extensive evaluations on the AIST++ dataset.
Figure 2: Exemplar videos generated from our DabFusion. Taking first image as starting frame and the unique music clip as guiding dance style, our framework is capable of generating varied styles of dance videos featuring different dancers from multiple perspectives with diverse initial poses and positions.
2 Related Works
Generating dance sequences from music uniquely intersects motion synthesis [1, 3, 11, 13] and music interpretation [57, 29, 58], aiming to create choreographed movements synchronized with input music. This extends beyond traditional motion synthesis, as choreographed movements are complex to animate. Early research [42, 68, 8] focused on producing 2D dance sequences due to the availability of online dance videos and advances in 2D human pose estimation [4]. However, 2D predictions lack expressiveness and applicability, prompting a shift to 3D dance generation. Recent methods leverage LSTMs [48, 19, 60], GANs [24, 47, 9, 20], and Transformers [16, 26, 46, 34, 63] for 3D motion generation. The AIST++ dataset [26], a rich compilation of 3D motion data, has significantly advanced the field. Notable models include FACT [26], a full-attention-based cross-modal transformer using sequence-to-sequence learning for lifelike 3D dance sequences, and Bailando [46], which combines a pose VQ-VAE with a Motion GPT for temporal coherence via actor-critic learning. While existing research focuses on generating motion sequences from music, Ren et al. [36] synthesizes videos using these sequences. Our work bypasses intermediary joint sequence generation, directly creating dance videos from music.
Image-to-Video (I2V) generation [2, 7, 15, 28, 33, 53, 59, 61, 65, 54] transforms static images into dynamic video sequences. This domain relates closely to single-image video prediction [21, 27], which infuses still images with temporal continuity to create believable motion. A specialized subfield, conditional image-to-video generation [2, 7, 15, 28, 53, 61, 54], incorporates guiding conditions or cues into the video generation process, enriching the content with specific characteristics or actions. For example, Wang et al. [53] developed a conditional GAN architecture to create human videos from static images, conditioned on labels describing facial expressions or actions. Yang et al. [61] introduced a pose-guided approach that extracts poses from an image and uses GANs to generate a sequence of poses and then video frames, translating static postures into animated sequences. Building on these, Wang et al. [54] unveiled the Latent Image Animator (LIA), which animates still images into videos under the guidance of an auxiliary video, leveraging a self-supervised auto-encoder to navigate latent space for animation. However, using audio, particularly music, as a conditioning factor in I2V remains largely unexplored. Introducing musical elements as conditions for video generation could open new research avenues and applications, suggesting an exciting direction for future work in I2V technology.
Figure 3: Overview of DabFusion. DabFusion incorporates noise input along with image embedding $z_{0}$ and music embedding $e$ as conditions. Following the denoising stage of the diffusion model, we obtain $h_{1}^{N}$, comprising a concatenated sequence of latent flow and corresponding occlusion maps. $h_{1}^{N}$ is utilized to transform $z_{0}$ into a new sequence of latent maps, denoted as $\tilde{z}_{1}^{N}$, which is subsequently decoded to produce an image sequence.
3 Method
Our methodology generates latent optical flows driven by musical inputs, with each music piece encoded by our specially designed music encoder. Music encoding is detailed in Sec. 3.1. Recent works [30, 43, 51, 52, 31] in motion transfer have shown the effectiveness of using latent optical flow for warping images. This approach is more resource-efficient, requiring less computational power and time compared to high-dimensional pixel or latent feature spaces. Our methodology begins by training an auto-encoder to discern optical flow between video frames, we discuss this in details in Sec. 3.2. This trained auto-encoder then aids in training the diffusion model to generate latent flows. The training of the diffusion model is detailed in Sec. 3.3. The overview of DabFusion is shown in Fig. 3 , given an initial image $x_{0}$ and a piece of music $m$, the image is first encoded into a latent space representation $z_{0}$, while the music is transformed into an embedding $e$. Subsequently, a volume of randomly sampled Gaussian noise, with dimensions $N\times H_{z}\times W_{z}\times 3$, undergoes a gradual denoising process by the U-Net. This process yields $h_{1}^{N}$ which is the concatenation of a latent flow sequence $f_{1}^{N}$ and corresponding occlusion map sequence $m_{1}^{N}$. $z_{0}$ is then warped by each latent flow in the latent flow sequence and corresponding occlusion map in the occlusion map sequence to generate a new latent map sequence $\tilde{z}_{1}^{N}$. These maps are sequentially fed into the image decoder, which synthesizes the frames of the new video.
3.1 Music Encoding
Extracting meaningful information from music is essential for generating dance videos, where dance type, dancer movement, and rhythm play critical roles. For dance type information, we utilize the CLAP model, a pipeline for contrastive language-audio pretraining that develops audio representations by integrating audio data with natural language descriptions. We fine-tune CLAP to enhance its capture of dance style information. Specifically, we freeze the weights of the audio and text encoders in CLAP and introduce multi-layer perceptrons as adapter layers. These layers are trained using data from the AIST++ dataset. For movement information, we employ Wav2CLIP, initially trained on audio-visual datasets. We apply a similar fine-tuning approach as with CLAP, using the AIST++ dataset to adapt Wav2CLIP for our specific requirements. Regarding rhythm, which is crucial for aligning dance movements with the music beat, we use Librosa, a tool for audio signal analysis, to extract beat information. We then concatenate these diverse sources of information to create a comprehensive musical representation tailored for dance video generation.
Figure 4: Training of latent flow auto-encoder. The flow predictor learns to estimate the latent flow $f$ and occlusion map $m$ between the reference frame $x_{ref}$ and the driving frame $x_{dri}$. The image encoder encodes $x_{ref}$ into a latent representation $z$, $f$ and $m$ are utilized to manipulate $z$ into $\tilde{z}$ which is then decoded by an image decoder to generate an output image $\hat{x}_{out}$. The objective of the training is to minimize the disparity between $x_{dri}$ and $\hat{x}_{out}$.
3.2 Latent Flow Estimation
The objective of this phase is the training of an auto-encoder capable of accurately capturing and modeling motion between video frames within a latent space. As illustrated in Fig. 4, the architectural components of the model comprise an image encoder, a flow predictor, and an image decoder. During the training process, two frames are randomly selected from a single video to serve as a reference frame $x_{ref}$ and a driving frame $x_{dri}$, both sharing dimensions of $H_{x}\times W_{x}\times 3$. The image encoder processes $x_{ref}$, converting it into a compact latent representation, $z$, with dimensions $H_{z}\times W_{z}\times C_{z}$. Subsequently, similar to [43, 52], the flow predictor, receiving both $x_{ref}$ and $x_{dri}$ as inputs, computes the backward latent optical flow, $f$, and an occlusion map, $m$, to denote the transformation between these frames. The estimated latent flow, $f$, maintains the spatial dimensions of $z$ and incorporates two channels to articulate the horizontal and vertical displacement across frames. Using backward flow estimation, the warping of $z$ using $f$ is realized through an efficient differentiable bilinear sampling [17]. Concurrently, the occlusion map, sized $H_{z}\times W_{z}\times 1$, facilitates the reconstruction of areas within $z$ that become obscured or revealed due to movement. Values within this map range from 0 to 1, where 1 signifies unoccluded regions, and 0 denotes complete occlusion. The warped latent map, $\tilde{z}$, is derived through the following equation:
$\tilde{z}\ =\ m\ \odot f_{w}(z\ ,\ f),$ (1)
where $f_{w}(\cdot\ ,\cdot)$ denotes the back-warping operation and $\odot$ is the element-wise multiplication. Finally, the image decoder takes $\tilde{z}$ as input, reconstructing visible portions while simultaneously inpainting occluded regions to produce the output image $\hat{x}_{out}$. The training objective is to minimize a reconstruction loss that quantifies the difference between $\hat{x}_{out}$ and $x_{dri}$. This loss employs the perceptual loss proposed by Johnson et al. [18], which utilizes features extracted by a pre-trained VGG-19 network [45]. Formally, the reconstruction loss is expressed as:
$\mathcal{L}_{rec}\ =\ \sum_{i=1}^{N}\left|V_{i}(\hat{x}_{out})-V_{i}(x_{dri})\right|,$ (2)
where $V_{i}(\cdot)$ extracts the $ith$ channel features from a specified layer of VGG-19, and $N$ is the total number of feature channels within that layer.
3.3 Latent Flow Generation
We use diffusion models to generate latent flow. These generative models reverse a diffusion process, which gradually turns data into Gaussian noise. The reverse process then restores this noise back to the original data distribution. Mathematically, the diffusion process is a Markov chain that adds noise to the data over $T$ steps. The process starts with the original data $x_{0}$ and ends with a sample $x_{T}$ that resembles Gaussian noise. The transition from $x_{t-1}$ to $x_{t}$ can be defined by the following equation:
$x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t},$ (3)
where $\alpha_{t}$ is a variance schedule that determines the amount of noise added at each step, and $\epsilon_{t}$ is sampled from a standard Gaussian distribution $\mathcal{N}(0,\mathcal{I})$. The reverse diffusion process seeks to reconstruct the original data from noise by learning the conditional distribution $p(x_{t-1}|x_{t})$. This is achieved by training a denoising model to estimate the parameters of the Gaussian distribution of $x_{t-1}$ given $x_{t}$. The model outputs parameters $\mu_{\theta}(x_{t},t)$ and $\sigma_{\theta}(x_{t},t)$, which define the Gaussian distribution from which $x_{t-1}$ is sampled:
$x_{t-1}=\mu_{\theta}(x_{t},t)+\sigma_{\theta}(x_{t},t)\cdot\epsilon,$ (4)
where $\epsilon$ is again sampled from a standard Gaussian distribution. The model is trained to minimize the difference between the generated and actual data distributions, with an objective function often approximating the negative log-likelihood of the data under the model. A practical approach for training diffusion models is employing the Mean Squared Error (MSE) between the predicted and actual noise used in data corruption at each diffusion step. Given the diffusion process, the goal during training is to predict the noise $\epsilon$ that was added to the original data at each step. The loss function for a single training step can be formulated as:
$\mathcal{L}=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}\right],$ (5)
where $\epsilon_{\theta}(x_{t},t)$ is the predicted noise by the model given the noised data $x_{t}$ at time step $t$. $\epsilon$ is the actual noise added to the data $x_{0}$ to obtain $x_{t}$.
Our approach utilizes a 3D U-Net [6] as the denoising mode. And the training of the model is in conjunction with the trained image encoder and flow predictor. Given an input video $x_{0}^{N}=\{x_{0},x_{1},...,x_{N}\}$ and its corresponding music $m$, we first use the flow predictor to estimate latent flow sequence $f_{1}^{N}=\{f_{1},...,f_{N}\}$ and occlusion map sequence $m_{1}^{N}=\{m_{1},...,m_{N}\}$ between starting frame $x_{0}$ and remaining frames $\{x_{1},...,x_{N}\}$. The size of $f_{1}^{N}$ is $N\times H_{z}\times W_{z}\times 2$ and the size of $m_{1}^{N}$ is $N\times H_{z}\times W_{z}\times 1$. Subsequently, we concatenate these two sequences along the last dimension to get $a_{0}=[f_{1}^{N},m_{1}^{N}]$ which has the size $N\times H_{z}\times W_{z}\times 3$. Then, we perform diffusion process which gradually adds 3D Gaussian noise to $a_{0}$ to map it to a a standard Gaussian noise. This is then reversed, using the image encoder to map the initial frame $x_{0}$ to latent space $z_{0}$ and incorporating music embeddings $e$ for conditioning the denoising model. The updated loss function, accounting for the conditioning on $z_{0}$ and $e$, is given by:
$\mathcal{L}=\mathbb{E}_{t,a_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(a_{t},t,z_{0},e)\|^{2}\right].$ (6)
Figure 5: Comparison of video quality between ground-truth videos and those generated by DabFusion. We use the same starting image and music piece to generate videos with our model, and select three frames from the same position.
4 Experiments and Results
4.1 Dataset and Implementation Details
Our model is trained and evaluated using the AIST++ [26] dataset, a leading resource in 3D human dance. AIST++ is recognized as the most comprehensive dataset in its domain, encompassing 10 dance genres. Each genre features 6 unique pieces of music, with each piece accompanying multiple videos, totaling 12,670 videos. These videos showcase a range of choreographies from basic to advanced, offering a wide spectrum of movements for detailed analysis. The pivotal advantage of AIST++ for our research is its meticulous synchronization of music and dance movements, which is crucial for music-driven dance generation and analysis.
We organize our dataset by music pieces, allocating 10,564 videos and 50 music tracks to the training set, and 2,106 videos with 10 tracks to the test set. Each video is segmented into frames, resized to $128\times 128$ pixels. In the latent flow estimation phase, we adopt the architecture from [18] for both our image encoder and decoder. The flow predictor follows [43]. The model is trained over 150 epochs with a batch size of 100, using the Adam optimizer [22], starting with a learning rate of $2\times 10^{-4}$, decreased by a factor of 0.1 after epochs 60, 90, and 120. For the denoising model’s training, we employ a conditional 3D U-Net architecture from [12], featuring four down-sampling and up-sampling 3D convolutional blocks. The time step $t$ is encoded by the sinusoidal position embedding [50]. The training spanned 250 epochs, starting with the same initial learning rate as the first stage, and adjusted at epochs 100, 150, and 200. We set $T$ = 1000 and apply a cosine noise schedule [32] and dynamic thresholding [40] at the 90% during sampling.
4.2 Evaluation Metrics
To assess the quality of generated videos, we use the Fréchet Video Distance (FVD) [49]. For quantitative assessment of image-level quality, we employ SSIM [55], PSNR [14], and LPIPS [66]. Additionally, we use the CLIP-Score (CS) [41] to measure the cosine similarity between CLIP embeddings of the dance video and Wav2CLIP embeddings of the music, evaluating the coherence between the dance video and the music. To evaluate synchronization between the dancer’s movements and the music beat, we introduce the 2D motion-music alignment score (2D-MM Align). This metric, inspired by the Beat Alignment Score [26] for 3D scenes, adapts the concept for 2D environments. The alignment score is calculated as the average distance between each kinematic beat and its closest music beat. 2D-MM Align is defined as:
[…]
where $F^{y}=\{f_{j}^{y}\}$ is the music beats, $F^{x}=\{f_{i}^{x}\}$ is the kinematic beats and $\sigma$ is a normalized parameter. Additionally, we incorporate the Audio-video Alignment Score (AV-Align) [62] to quantify audio-video synchronization. This metric analyzes optical flow in the video stream and audio peaks in the audio stream and assesses their temporal alignment within a three-frame window.
Figure 6: Exemplar videos generated using our DabFusion, showcasing dance videos with unseen individuals in various poses.
Table 1: Quantitative results between DabFusion and MM-Diffusion.
[…]
MM-Diffusion 180.31 0.025 34.81 0.951
DabFusion (ours) 193.98 0.033 26.88 0.959
4.3 Result Analysis
Video Quality Evaluation. To evaluate the quality of the videos generated by our DabFusion, we use the music and the first frame from the test set, resulting in 2,106 generated videos. We visually present outputs from our model alongside ground truth in Fig. 5. As shown in Fig. 5, the quality of the videos generated by our model closely approximates that of real videos. Given the novelty of our task and the lack of directly comparable previous works, we benchmark our DabFusion against MM-Diffusion [39], the leading model in unconditional video generation. We select MM-Diffusion because it is trained on a subset of the AIST++ dataset, closely aligning with our training dataset and demonstrating state-of-the-art video quality performance on AIST++. We randomly generate an equivalent number of videos with MM-Diffusion for comparison. A quantitative comparison between our models and MM-Diffusion is provided in Tab. 1. The results reveal that the quality of videos produced by our models closely matches that of MM-Diffusion. These comparisons underscore the competitive quality of videos generated by our approach.
Table 2: Quantitative results across different angles. C01 is in front of the dancer, C02 is in the upper right of the dancer, C03 is in the right of the dancer, C09 is in front of the dancer but closer than C01.
[…]
MM-Diffusion C01 176.93 0.019 35.52 0.949
DabFusion (ours) C01 158.02 0.027 26.33 0.964
C02 198.17 0.040 25.23 0.954
C03 157.94 0.036 26.29 0.960
C09 195.17 0.035 25.28 0.956
A notable characteristic of the AIST++ dance dataset is its use of nine cameras C01-C09 to capture dance performances, offering multiple perspectives. We analyze four specific angles, summarized in Tab. 2, to assess their impact on video quality. MM-Diffusion, trained on a subset of the AIST++ dataset using only data from the C01 camera, is used to generate an equivalent number of videos for direct comparison. Our models outperform MM-Diffusion in terms of FVD and SSIM scores, especially for camera position C01, demonstrating superior video quality. Analysis of scores among different cameras indicates that direct camera angles, such as C01 and C03, where the cameras point straight at the dancer, result in higher-quality video generation compared to oblique angles like C02. Additionally, we explore the impact of camera distance by comparing videos from C01 and C09, which share the same angle but differ in their proximity to the dancer. The findings suggest that cameras positioned further from the dancer tend to yield better results.
Table 3: Alignment results for DabFusion.
[…]
Ground Truth 0.153 0.223 0.152
DabFusion (ours) 0.142 0.215 0.148
Table 4: Alignment results based on dance styles.
[…]
Break Ground Truth 0.138 0.231 0.226
DabFusion (ours) 0.131 0.223 0.229
Ballet Jazz Ground Truth 0.063 0.198 0.101
DabFusion (ours) 0.057 0.193 0.098
House Ground Truth 0.121 0.218 0.134
DabFusion (ours) 0.112 0.213 0.128
Alignment Evaluation. In Tab. 3, we assess the synchronicity between music-video and music-motion pairings. The results indicate that videos generated by our DabFusion closely match the ground-truth videos in terms of alignment metrics. We also evaluate the results according to dance style categories corresponding to the music. These assessments, detailed in Tab. 4, exhibit significant disparities in alignment scores among various dance genres. Notably, Ballet Jazz yields the lowest alignment scores, attributed to its challenging choreography compared to other dance styles. Consequently, the alignment scores reflect the varying levels of difficulty inherent to each dance genre.
4.4 Choreograph Anyone
DabFusion can choreograph individuals whether they belong to the AIST++ dataset or not, requiring several preprocessing steps. First, we use YOLO [35] for object detection and the Segment Anything Model [23] for segmentation to identify and segment the person from unseen scenarios. We then substitute the individual in the image from the AIST++ dataset with the newly segmented person, creating a combined image. This fused image is set as the starting frame to generate dance videos with musical conditions. We randomly select images featuring diverse individuals in various poses to create dance videos. As illustrated in Fig. 6, DabFusion is capable of generating high-quality dance videos featuring previously unseen individuals in a variety of contexts and poses.
Table 5: Ablation study for result comparison between 40-frame video and 80-frame video.
[…]
40 frames 0.142 0.215 0.148 193.98
80 frames 0.153 0.219 0.179 194.84
Table 6: Ablation study for result comparison between DabFusion with and without beat information.
[…]
DabFusion $\times$ 0.138 0.202 0.141 194.44
$\checkmark$ 0.142 0.215 0.148 193.98
Table 7: Ablation study for result comparison about arbitrary length video generation. We use the last frame of the first generated videos as the starting frame of next video to generate longer videos.
[…]
40 frames 0.142 0.215 0.148 193.98
80 frames 0.145 0.203 0.139 210.86
120 frames 0.131 0.197 0.121 256.32
160 frames 0.123 0.179 0.101 323.72
4.5 Ablation Study
Influence of Video Length. We observe that video length impacts benchmark performance. Our model is trained to generate videos of 40 and 80 frames. Due to training time and computational resource considerations, we designate the 40-frame videos as the baseline. This comparison is detailed in Tab. 5. The results indicate comparable video quality between the two lengths; however, the 80-frame videos demonstrate superior performance in alignment evaluations. This improvement is primarily attributed to the extended music duration, providing more information for enhancing alignment accuracy.
Influence of Beat Information. With the emergence of large foundation models capable of generating music representation, many current approaches rely exclusively on these models to encode music, often overlooking information from traditional audio analysis tools. In our experiment, we explore the impact of incorporating beat information into our dance video generation process. The findings, presented in Tab. 6, show that while the inclusion of beat information does not substantially alter the overall quality of the generated videos, it does enhance the performance in terms of video-audio synchronization and motion-beat alignment. Specifically, the model that integrates beat information outperforms the one without in all assessed metrics, underscoring the beneficial effects of beat information.
Arbitrary-Length Video Generation. Given the specialty of our task, our model theoretically can generate videos corresponding to any length of music. By using the final frame of one video sequence as the initial frame of the subsequent sequence, we can create videos of arbitrary length. We conduct experiments with four varying lengths of videos generated sequentially using our model; the results are presented in Tab. 7. As video length increases, quality tends to drop rapidly. This decline is primarily because the final frames generated by the model lack the fidelity of actual video frames, becoming more apparent with longer videos. Additionally, alignment scores diminish due to this quality degradation. For instance, employing a pose estimation algorithm to determine keypoints for calculating the 2D-MM Alignment score becomes less accurate with lower video quality, adversely affecting the alignment score.
5 Conclusion
We introduce the challenge of creating dance videos from music with our model, DabFusion. This task eliminates the need for keypoint annotations, simplifying data use and enhancing dataset compatibility. DabFusion can animate individuals from reference images, aligning their movements to various musical styles. Our specially designed music encoder effectively captures essential features including dance style, movement, and rhythm, vital for dynamic choreography. Using the AIST++ dataset, we demonstrate DabFusion’s ability to produce high-quality dance videos with excellent synchronization between audio and video, setting a strong foundation for further research in dance video generation.